5. Determine the Suppose Y has the binomial pmf. p PY-1)- 40 and n standard deviation of Y. Express your answer to three decimal places
Answer #1

Here we have

p=.40 and n=5

So standard deviation(\sigma)=sqrt(p*(1-p)*n)=sqrt(.4*.6*5)=1.095